OpenStudy (anonymous):

Evaluate the expression for the given values of the variables. 4t + 2u2 – u3; t = 2 and u = 1

OpenStudy (goformit100):

PUT THE VALUES IN IT DIRECTLY

OpenStudy (anonymous):

3/5 = x /3 CROSS MULTIPLY 5x = 9 x = 9/5 OR 1.8 ANSWER 2) 4*2 + 2*1^2 - 1^3 = 8 +2 -1 = 9 ANSWER.

OpenStudy (anonymous):

i think

OpenStudy (goformit100):

WHERE IS X HERE?

OpenStudy (anonymous):

9

OpenStudy (goformit100):

YES YOU ARE CORRECT

OpenStudy (anonymous):

:)